solve y(2x+y-2)dx=2(x+y)dy


1
Expert's answer
2021-12-21T04:13:02-0500

Solution:

On comparing given equation with "M(x,y)dx + N(x,y)dy =0" , we get,

"M(x,y) = y(y + 2x -2 ) , N(x,y) = -2(x + y)"

The equation is not exact because

"M_y =2(x + y -1) ; N_x = -2." But "( M_y - N_x )\/N =- 1."

So the "I.F. = e^{-x}"

may be used to obtain the exact equation "P(x,y)dx + Q(x,y)dy = 0"

with

"P(x,y) = (e^{-x})(y(y+2x - 2)) , Q(x,y) = - 2e^{-x}(x+y)" and "P_y = Q_x \n\n= e^{-x} 2( y+x-1)." The equation is an exact differential dF(x,y) = 0 then, the solution

is F(x,y) = C, with "F_x =P = (e^{-x})( y( y +2x - 2))" ...(1) ,

"F_y = Q = - 2e^{-x}(x + y)" ...(2).

Integrating eq. (1) yields "F(x,y) = -e^{-x} y( y + 2x) + G(x)" and using eq.(2) gives

G’(x) =0 so G(x) =C .

Then the solution is : "F(x,y) = - e^{-x} y (y+2x) = C"